OpenStudy (anonymous):

Create the division of two monomials by filling in the blanks with exponents greater than one. Find the quotient of these monomials. Explain which operation was done on the exponents to find this quotient. Prove this operation provides the correct quotient by showing the fraction in expanded form.

OpenStudy (anonymous):

a6/a3

OpenStudy (jim766):

do you know how to divide this?

OpenStudy (jim766):

a^6/a^3 = a^3

OpenStudy (jim766):

expanded form \[a^6/a^3 = a*a*a*a*a*a/a*a*a\]